1,400 tons of wheat of two different varieties was delivered to the silo. When processing one variety, there was 2% of waste and 3% during the processing of the second variety of wheat. After the processing, the amount of remaining wheat was 1,364 tons. How many tons of each variety of wheat was delivered to the silo?

Answer by ikleyn(53763)

.
1,400 tons of wheat of two different varieties was delivered to the silo. When processing one variety, there was 2% of waste
and 3% during the processing of the second variety of wheat. After the processing, the amount of remaining wheat was 1,364 tons.
How many tons of each variety of wheat was delivered to the silo?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~

    x +     y = 1400          (1)
0.02x + 0.03y = 1400-1364.    (2)

Simplify:

    x +     y = 1400          (1')
0.02x + 0.03y =   36.         (2')

Multiply the first eq. by 2. Multiply the second eq. by 100. You will get

2x +  y = 2800,               (1'')
2x + 3y = 3600.               (2'')

Subtract eq. ('1'') from the eq. (2''). You will get

2y = 800.

From this point please complete the solution on your own.

If you have questions or want to say "thanks", then please mention the ID number of this problem (# 1065528) in your response.